# Benchmarking Pre-trained Large Language Models’ Potential Across Urdu NLP tasks

### v1

@@ 4 Conclusion and Future Work
In this study, we benchmark the potential of both open and closed LLMs on 14 Urdu NLP tasks employing a substantial number of publicly accessible datasets. Through our experiments we provide a comparative performance analysis for each task and dataset against the SOTA. These findings will assist the Urdu NLP community in selecting suitable models for usage and fine-tuning within specific contexts. As future work, we aim to develop a public leaderboard for Urdu benchmarking and explore integration of additional models, tasks, and datasets. We further aim to refine prompt engineering, investigate few-shot and fine-tuning settings to minimize the performance gap with the SOTA.
@@
Our study is confined to three LLMs and does not include the heavier versions of models such as Bloomz-170B or Llama 2 70B due to hardware and computational resource limitations which may impact the comprehensiveness of the analysis. This limitation may affect the generalizability of the findings to models with higher parameters, potentially missing insights into the performance of more robust versions of these language models. Our study also primarily concentrates on evaluating the models in a zero-shot setting. While this setting provides valuable insights into the models’ out-of-the-box performance, it may not capture the full potential of fine-tuned models for specific tasks. Our study also does not extensively delve into the quality and representativeness of the training data for Urdu language used in these models.
@@ Appendix A Appendix
A.1 Prompts - Named Entity Recognition
@@ A.1.3 Llama 2
<<SYS>> You are Performing Name Entity Recognition for the urdu words.<</SYS>> Human: Word: Please select one of the following entity: Person Organization Location Other No explanation or further assistance is required. Only entity name is required Assistant: The entity is
@@ A.8.1 Bloomz
system: You are an expert in detecting threat in the urdu samples
Classify the sentence as threatening or non threatening. Output "1" for threatening and "0" for non threatening. Sentence: :
@@ A.9.3 Llama 2
<<SYS>> You are a helpful assistant in classification of cyberbullying. You should always provide answer from given labels without explanation. <</SYS>> Human: Sentence . classify the nature of cyber bullying present in sentence with one of the following label: INSULT OFFENSIVE NAMECALLING PROFANE THREAT CURSE NONE Assitant:
@@ A.11.3 Llama 2
A.12 Prompts - Text Summarization
@@ A.12.3 Llama 2
A.13 Prompts - Sentiment Analysis
@@ A.13.1 Bloomz
A.13.2 GPT 3.5
"system": "You are an expert in detecting abusive language in the urdu samples” Do the sentiment analysis. Output should be "pos" for positive sentences , "neu" for neutral sentences and "neg" for negative sentences. No explanation is required. Sentence:
A.13.3 Llama 2
<<SYS>> You are a helpful assistant in sentiment analysis. You should always provide answer from given labels without explanation. <</SYS>> Human: Do the sentiment analysis. Output "neu" for neutral sentence, "pos" for positive sentence , and "neg" for negative sentence. No explanation is required. Sentence: Assistant:
A.14 Prompts - Sentiment Analysis (CLE)
A.14.1 Bloomz
@@
A.14.3 Llama 2
<<SYS>> You are a helpful assistant in sentiment analysis. You should always provide answer from given labels without explanation. <</SYS>> Human: Tweet: . Perform sentiment analaysis on the tweet and answer with one of the following label: -2 : Highly negative -1 : Negative 0 : Neutral 1 : Positive 2 : Highly positive Assitant:
A.15 Prompts - Multi-label Emotion Classification
A.15.1 Bloomz
@@ A.15.3 Llama 2
A.16 Prompts - Emotion Classification
A.16.1 Bloomz
@@ A.16.3 Llama 2
A.17 Prompts - Machine Translation
A.17.1 Bloomz
@@
A.17.2 GPT 3.5
"system": "You are an expert translator specialized in translating texts from English to Urdu "
Translate the following English sentence to Urdu:
A.17.3 Llama 2

### v2

@@ 4 Conclusion and Future Work
In this study, we benchmark the potential of both open and closed LLMs on 17 Urdu NLP tasks employing a substantial number of publicly accessible datasets. Through our experiments we provide a comparative performance analysis for each task and dataset against the SOTA. These findings will assist the Urdu NLP community in selecting suitable models for usage and fine-tuning within specific contexts. As future work, we aim to develop a public leader board for Urdu benchmarking and explore integration of additional models, tasks, and datasets. Also, after evaluating multiple models, we are focusing on pretraining Llama 3.1 to enhance Urdu language support and expand token coverage for greater adaptability across diverse languages and domains. This will also includes domain-specific fine-tuning to further boost performance, leveraging Llama 3.1’s compact size, active community, and robust results.
@@
Our study is confined to seven LLMs and does not include the heavier versions of models such as Bloomz-170B or Llama 3.1 405B due to hardware and computational resource limitations which may impact the comprehensiveness of the analysis. This limitation may affect the generalization of the findings to models with higher parameters, potentially missing insights into the performance of more robust versions of these language models. Our study also primarily concentrates on evaluating the models in a zero-shot setting. While this setting provides valuable insights into the models’ out-of-the-box performance, it may not capture the full potential of fine-tuned models for specific tasks. Our study also does not extensively delve into the quality and representativeness of the training data for Urdu language used in these models.
Acknowledgments
We would like to express our sincere gratitude to Dr. Miriam Butt for allowing us to utilize the infrastructure at the University of Konstanz for our benchmarking experiments. Additionally, we extend our acknowledgment to Miss Ayesha Khalid for supplying the ASR dataset and corresponding SOTA techniques.
@@ Appendix A Appendix
A.1 Prompts - Name Entity Recognition
@@ A.1.3 Llama 2
You are Performing Name Entity Recognition for the urdu words.<</SYS>> Human: Word: Please select one of the following entity: Person Organization Location Other No explanation or further assistance is required. Only entity name is required Assistant: The entity is
A.1.4 Llama 3.1
You are a name entity recognition model. Your task is to mark the entity as Person, Organization, Location, or Other in Urdu text samples. Ensure that your outputs are Person, Organization, Location, or Other. No explanation is required.
A.1.5 Ministral 8B
@@ A.3.3 Llama 2
A.3.4 Llama 3.1
You are an intent classification model. Your task is to mark the intent as 0 or 1 or 2 in Urdu text samples. Ensure that the model outputs ’0’ for Informational intent , ’1’ for Navigational intent and ’2’ for Transitional intent. Ensure that your outputs 0 or 1 or 2 only. No explanation is required.
A.3.5 Ministral 8B
You are an intent classification model. Your task is to identify the intent in the following urdu sentence. Intents are: Informational Navigational Transitional Only output the name of the intent. No explanation is required.
@@ A.4.3 Llama 2
A.4.4 Llama 3.1
You are a hate speech classification model. Your task is to mark 0 or 1 or 2 or 3 in Urdu text samples. Ensure that the model outputs ’1’ for Sectarian hate text , ’2’ for Interfaith hate text , ’3’ for Ethnic hate text and ’0’ for any other hate text. Ensure that your outputs 0 or 1 or 2 or 3 only. No explanation is required.
A.4.5 Ministral 8B
You are a hate speech classification model. Labels: 1: Sectarian hate 2: Interfaith hate 3: Ethnic hate 0: None of the above Instructions: To distinguish between hate speech and non-hate speech in text samples. Ensure that the model output "1" for hate related to "Sectarian", "2" for hate related to "Interfaith" and "3" for hate related to "Ethnic" and "0" if you think it does not fall in these three categories. Your output should be only 0, 1, 2 or 3. No explanation is required. Sentence: Label(0,1,2,3):
@@ A.5.3 Llama 2
A.5.4 Llama 3.1
You are a hate speech classification model. Your task is to mark 0 or 1 in Urdu text samples. Ensure that the model outputs ’0’ for Non-hate text and ’1’ for Hate text. Ensure that your outputs 0 or 1 only. No explanation is required.
A.5.5 Ministral 8B
@@ A.7.3 Llama 2
A.7.4 Llama 3.1
You are a abusive language detection model.
Labels:
0: non-abusive language 1: abusive language Instructions: To distinguish between abusive and non-abusive language in text samples. Ensure that the model outputs 0 for non-abusive language and 1 for abusive language. Your output should be only 0 or 1. No explanation Sentence: Label(0 or 1):
A.7.5 Ministral 8B
@@ A.8.1 Bloomz
system: You are an expert in detecting threat in the urdu samples Classify the sentence as threatening or non threatening. Output "1" for threatening and "0" for non threatening. Sentence: :
@@ A.8.3 Llama 2
A.8.4 Llama 3.1
You are a classification model. Your job is to classify the sentences as threatening or non-threatening. Output "1" if the sentence is threatening and "0" if it is non-threatening. No explanation is required
A.8.5 Ministral 8B
Classify the sentence as threatening or non threatening. Output "1" for threatening and "0" for non threatening. No explanation is required. Sentence:
@@ A.9.3 Llama 2
You are a helpful assistant in classification of cyberbullying. You should always provide answer from given labels without explanation. <</SYS>> Human: Sentence . classify the nature of cyber bullying present in sentence with one of the following label: INSULT OFFENSIVE NAMECALLING PROFANE THREAT CURSE NONE Assitant:
A.9.4 Llama 3.1
You are a cyberbullying classification model. Your task is to mark the input as INSULT or OFFENSIVE or NAMECALLING or PROFANE or THREAT or CURSE or NONE in Urdu text samples. Ensure that your output is INSULT or OFFENSIVE or NAMECALLING or PROFANE or THREAT or CURSE or NONE. No explanation is required.
A.9.5 Ministral 8B
@@ A.10.3 Llama 2
A.10.4 Llama 3.1
You are a fake news detection model.Output "fake" for fake news and "real" for real news. No explanation is required.
A.10.5 Ministral 8B
@@ A.11.3 Llama 2
A.11.4 Llama 3.1
You are a hate speech classification model. Your task is to mark 0 or 1 in Urdu text samples. Ensure that the model outputs ’0’ for Non-hate text and ’1’ for Hate text. Ensure that your outputs 0 or 1 only. No explanation is required.
A.11.5 Ministral 8B
You are a hate speech classification model. Your task is to mark 1 or -1 in Urdu text samples. Ensure that the model outputs ’1’ for Non-hate text and ’-1’ for hate text. Ensure that your outputs 1 or -1 only. No explanation is required.
A.12 Prompts - Extractive Summarization
@@ A.12.3 Llama 2
A.12.4 Llama 3.1
You are an extractive summarization model. You have to decide whether the given passage should be kept or discarded for summary? Output "1" if the passage should be kept and "0" for discarding (0 or 1) without explanation
A.12.5 Ministral 8B
You are an extractive summarization model. You have to decide whether the given passage should be kept or discarded for summary? Output "1" if the passage should be kept and "0" for discarding (0 or 1) without explanation
A.13 Prompts - Abstractive Summarization
@@ A.13.1 Bloomz
Write summary of the given Urdu meeting. Meeting: . Summary:
A.13.2 GPT 3.5
You are a summarization model. Generate summary for the given meeting minutes in Urdu. Prompt: Meeting minutes: Summary:
A.13.3 Llama 2
You are a summarization model. Your job is to summarize the urdu meeting minutes. Meeting minutes: Summary:
A.13.4 Llama 3.1
You are a summarization model. Generate the summary for the meeting minutes in Urdu.
A.13.5 Ministral 8B
You are a summarization model. Generate the summary for the meeting minutes in Urdu.
A.14 Prompts - Sentiment Analysis
A.14.1 Bloomz
@@
A.14.2 GPT 3.5
Do the sentiment analysis. Output should be "pos" for positive sentences , "neu" for neutral sentences and "neg" for negative sentences. No explanation is required. Sentence:
A.14.3 Llama 2
You are a helpful assistant in sentiment analysis. You should always provide answer from given labels without explanation. Human: Do the sentiment analysis. Output "neu" for neutral sentence, "pos" for positive sentence , and "neg" for negative sentence. No explanation is required. Sentence: Assistant:
A.14.4 Llama 3.1
Perform sentiment analysis. Your output should be "pos" for positive sentence , "neu" for neutral sentence and "neg" for negative sentence. No explanation is required.
A.14.5 Ministral 8B
You are a helpful assistant in sentiment analysis. You should always provide answer from given labels without explanation. Tweet: . Perform sentiment analaysis on the tweet and answer with one of the following label: -2 for Highly negative -1 for Negative 0 for Neutral 1 for Positive 2 for Highly positive No explanation is required. Output (-2,-1,0,1,2):
A.15 Prompts - Sentiment Analysis (CLE)
A.15.1 Bloomz
@@
A.15.3 Llama 2
You are a helpful assistant in sentiment analysis. You should always provide answer from given labels without explanation.
Human: Tweet: . Perform sentiment analysis on the tweet and answer with one of the following label: -2 : Highly negative -1 : Negative 0 : Neutral 1 : Positive 2 : Highly positive Assitant:
A.15.4 Llama 3.1
You are a helpful assistant in performing sentiment analysis. Perform sentiment analaysis on the tweet and answer with one of the following label: -2 for Highly negative -1 for Negative 0 for Neutral 1 for Positive 2 for Highly positive Only output -2 to 2 based on the above mentioned scale. No explanation is required
A.15.5 Ministral 8B
You are a helpful assistant in performing sentiment analysis. Perform sentiment analaysis on the tweet and answer with one of the following label: -2 for Highly negative -1 for Negative 0 for Neutral 1 for Positive 2 for Highly positive Only output -2 to 2 based on the above mentioned scale. No explanation is required
A.16 Prompts - Multi-label Emotion Classification
A.16.1 Bloomz
@@ A.16.3 Llama 2
A.16.4 Llama 3.1
Output the emotion or emotions(if multiple) for the sentence. Emotions: anger, disgust, fear, sadness, surprise, happiness, neutral. You can output multiple emotions as well but should only be the name of the emotions. No explanation is required.
A.16.5 Ministral 8B
@@ A.17.3 Llama 2
A.17.4 Llama 3.1
Output the emotions for the paragraph from one of the following: neutral, happy, fear, sad, anger, love. Your output should only be the name of one of the given emotions. Don’t provide any other apart from these six emotions. No explanation is required
A.17.5 Ministral 8B
Output the emotions for the paragraph from one of the following: neutral, happy, fear, sad, anger, love. Your output should only be the name of one of the given emotions. Don’t provide any other apart from these six emotions. No explanation is required
A.18 Prompts - Machine Translation
A.18.1 Bloomz
@@
A.18.2 GPT 3.5
"system": "You are an expert translator specialized in translating texts from English to Urdu " Translate the following English sentence to Urdu:
A.18.3 Llama 2
@@
A.18.4 Llama 3.1
You are an English to Urdu translator. Translate the english sentences into Urdu. No explanation is required. Just translate into Urdu
A.18.5 Ministral 8B
You are an expert translator specialized in translating texts from English to Urdu. Translate the following English sentence to Urdu: "". Provide only the Urdu translation, without any additional text or explanations.
A.19 Prompts - POS Tagging
A.19.1 Bloomz
Your task is to tag POS in input. You will use following Taggig scheme: Tag Proper Noun as NNP ,Tag Common Noun as NN,Tag Personal pronoun as PRP,Tag Demonstrative as PDM,Tag Possessive pronouns as PRS,Tag Reflexive pronouns as PRF,Tag Reflexive Apna as APNA,Tag Relative Personal as PRR,Tag Relative Demonstrative as PRD,Tag Main Verb Infinitive as VBI,Tag Main Verb Finite as VB,Tag Aspectual auxiliaries as AUXA,Tag Progressive auxiliaries as AUXP,Tag Tense auxiliaries as AUXT,Tag Modals auxiliaries as AUXM,Tag Foreign Fragment as FF,Tag Interjection as INJ,Tag Preposition as PRE,Tag Postposition as PSP,Tag Common as SYM,Tag Punctuation as PU,Tag Common as RB,Tag Negation as NEG,Tag Common as PRT,Tag Vala as VALA,Tag Coordinate Conjunction as CC,Tag Subordinate Conjunction as SC,Tag SC Kar as SCK,Tag Pre-sentential as SCP,Tag Ordinal as OD,Tag Fraction as FR,Tag Multiplicative as QM,Tag Adjective as JJ,Tag Quantifier as Q,Tag Cardinal as CD. Your Output should be only one tag corrosponding to input word. no explaination is required. input:
A.19.2 GPT 3.5
"system": "You are an expert in Urdu pos tagging " Your task is to tag POS in input. You will use following Taggig scheme: Tag Proper Noun as NNP ,Tag Common Noun as NN,Tag Personal pronoun as PRP,Tag Demonstrative as PDM,Tag Possessive pronouns as PRS,Tag Reflexive pronouns as PRF,Tag Reflexive Apna as APNA,Tag Relative Personal as PRR,Tag Relative Demonstrative as PRD,Tag Main Verb Infinitive as VBI,Tag Main Verb Finite as VB,Tag Aspectual auxiliaries as AUXA,Tag Progressive auxiliaries as AUXP,Tag Tense auxiliaries as AUXT,Tag Modals auxiliaries as AUXM,Tag Foreign Fragment as FF,Tag Interjection as INJ,Tag Preposition as PRE,Tag Postposition as PSP,Tag Common as SYM,Tag Punctuation as PU,Tag Common as RB,Tag Negation as NEG,Tag Common as PRT,Tag Vala as VALA,Tag Coordinate Conjunction as CC,Tag Subordinate Conjunction as SC,Tag SC Kar as SCK,Tag Pre-sentential as SCP,Tag Ordinal as OD,Tag Fraction as FR,Tag Multiplicative as QM,Tag Adjective as JJ,Tag Quantifier as Q,Tag Cardinal as CD. Your Output should be only one tag corrosponding to input word. no explaination is required. input:
A.19.3 Llama 2
Your task is to tag POS in input. You will use following Taggig scheme: Tag Proper Noun as NNP ,Tag Common Noun as NN,Tag Personal pronoun as PRP,Tag Demonstrative as PDM,Tag Possessive pronouns as PRS,Tag Reflexive pronouns as PRF,Tag Reflexive Apna as APNA,Tag Relative Personal as PRR,Tag Relative Demonstrative as PRD,Tag Main Verb Infinitive as VBI,Tag Main Verb Finite as VB,Tag Aspectual auxiliaries as AUXA,Tag Progressive auxiliaries as AUXP,Tag Tense auxiliaries as AUXT,Tag Modals auxiliaries as AUXM,Tag Foreign Fragment as FF,Tag Interjection as INJ,Tag Preposition as PRE,Tag Postposition as PSP,Tag Common as SYM,Tag Punctuation as PU,Tag Common as RB,Tag Negation as NEG,Tag Common as PRT,Tag Vala as VALA,Tag Coordinate Conjunction as CC,Tag Subordinate Conjunction as SC,Tag SC Kar as SCK,Tag Pre-sentential as SCP,Tag Ordinal as OD,Tag Fraction as FR,Tag Multiplicative as QM,Tag Adjective as JJ,Tag Quantifier as Q,Tag Cardinal as CD. Your Output should be only one tag corrosponding to input word. no explaination is required. input:
A.19.4 Llama 3.1
Your task is to tag POS in input. You will use following Taggig scheme: Tag Proper Noun as NNP ,Tag Common Noun as NN,Tag Personal pronoun as PRP,Tag Demonstrative as PDM,Tag Possessive pronouns as PRS,Tag Reflexive pronouns as PRF,Tag Reflexive Apna as APNA,Tag Relative Personal as PRR,Tag Relative Demonstrative as PRD,Tag Main Verb Infinitive as VBI,Tag Main Verb Finite as VB,Tag Aspectual auxiliaries as AUXA,Tag Progressive auxiliaries as AUXP,Tag Tense auxiliaries as AUXT,Tag Modals auxiliaries as AUXM,Tag Foreign Fragment as FF,Tag Interjection as INJ,Tag Preposition as PRE,Tag Postposition as PSP,Tag Common as SYM,Tag Punctuation as PU,Tag Common as RB,Tag Negation as NEG,Tag Common as PRT,Tag Vala as VALA,Tag Coordinate Conjunction as CC,Tag Subordinate Conjunction as SC,Tag SC Kar as SCK,Tag Pre-sentential as SCP,Tag Ordinal as OD,Tag Fraction as FR,Tag Multiplicative as QM,Tag Adjective as JJ,Tag Quantifier as Q,Tag Cardinal as CD. Your Output should be only one tag corrosponding to input word. no explaination is required. input:
A.19.5 Ministral 8B
Your task is to tag POS in input. You will use following Taggig scheme: Tag Proper Noun as NNP ,Tag Common Noun as NN,Tag Personal pronoun as PRP,Tag Demonstrative as PDM,Tag Possessive pronouns as PRS,Tag Reflexive pronouns as PRF,Tag Reflexive Apna as APNA,Tag Relative Personal as PRR,Tag Relative Demonstrative as PRD,Tag Main Verb Infinitive as VBI,Tag Main Verb Finite as VB,Tag Aspectual auxiliaries as AUXA,Tag Progressive auxiliaries as AUXP,Tag Tense auxiliaries as AUXT,Tag Modals auxiliaries as AUXM,Tag Foreign Fragment as FF,Tag Interjection as INJ,Tag Preposition as PRE,Tag Postposition as PSP,Tag Common as SYM,Tag Punctuation as PU,Tag Common as RB,Tag Negation as NEG,Tag Common as PRT,Tag Vala as VALA,Tag Coordinate Conjunction as CC,Tag Subordinate Conjunction as SC,Tag SC Kar as SCK,Tag Pre-sentential as SCP,Tag Ordinal as OD,Tag Fraction as FR,Tag Multiplicative as QM,Tag Adjective as JJ,Tag Quantifier as Q,Tag Cardinal as CD. Your Output should be only one tag corrosponding to input word. no explaination is required. input: